- Remote participants by eZuce: None
Report of work (*:responsible person)
- VAC: *Uchiyama, Kimura, Suzuki
- Pump down started yesterday, but a big leak was found. We changed a flange at BS and it is ready to pump down again today.
- Alignment, ETMY: *Enomoto, Miyo, Miyamoto, Izumi
- Yaw damping stared working by changing polarity.
- Beam alignment to both end and REFL port was confirmed the last night.
- ~10rad shift was seen as usual.
- Clipping was fixed by PSL alignment yesterday.
- PR3 is being shifted to yaw by pump down by ~300urad.
- LSC model as full model started running at k1imc0. Dolhpin will be conected soon.
- Camera: Yokozawa
- IYAC and IXAC areas are being pumped down.
- CRY: Mirror reached at 18K.
Work plan (*:responsible person)
- VAC: Uchiyama
- Pump down.
- DGS: *Miyakawa
- 10GB network card at dc0 will be replaced.
- [AM] IR: PSL, Green: off.
- [PM] IR: PSL, Green: off.
- [night] IR: all, Green: off
- week of 3/12
- TMS installation at ETMY.
- Connecting ducts at the corner area.
- week of 3/19
- Vacuum test at at the corner area.
- TMS alignment using green light at ETMY.
- Beam alignment to REFL port.
- Connection of PD signal to DAQ should be implemented this week or next week.